OpenStudy (anonymous):

Find the equation of the circle with center at (3, 2) and through the point (5, 4).

OpenStudy (ranga):

Find the distance between the two points and that will be the radius. Then: (x-h)^2 + (y-k)^2 = r^2 where (h,k) is the center and r is the radius.

OpenStudy (anonymous):

so it would be (x-3)^2 + (y-2)^2 = 2^2

OpenStudy (ranga):

on the right I am getting 8 because radius is sqrt(8).

OpenStudy (anonymous):

ok gotcha so (x-3)^2 + (y-2)^2 = 8?

OpenStudy (ranga):

yes. radius r = sqrt( (4-2)^2 + (5-3)^2 ) = sqrt(8)